OpenStudy (anonymous):

Please help, I don't understand this question. 2. Suppose that y varies directly with x and inversely with z, and y = 18 when x = 15 and z = 5. Write the equation that models the relationship. Then find y when x = 21 and z = 7. (1 point) y = 5z /x; 5/ 3 y = 5x/z; 15 y = 6x/z; 18 y = 6z/x; 2

hero (hero):

Hint: yz = kx

hero (hero):

1. Plug the 1st set of given values into the formula 2. Solve for k 3. Plug the second set of given values into the formula including the value for k 4. Solve for y

hero (hero):

`1.` (18)(5) = 15k 90 = 15k 90/15 = k `2.` 6 = k `3.` 7y = (6)(21) 7y = 126 y = 126/7 `4.` y = 18

OpenStudy (anonymous):

Thank you so much. I just got back from class.